A sample of helium has a temperature of 120. °C and pressure of 746 Torr. What is the temperature (in °C), when the pressure is 715 Torr?
Answer #1

Ideal cas &qn is :- [PV = nRT] where P= Pressure N-Volume menol, R- Las Constant , S T = Temperatune. P = Pz → 746 715 (120+2一步 | Tv = 315 ¥32= = 334.6k HE Timec = 216.61-233 “T.= 103.61]